# mGTE: Generalized Long-Context Text Representation and Reranking Models for Multilingual Text Retrieval

### v1

@@ Abstract
We present systematic efforts in building long-context multilingual text representation model (TRM) and reranker from scratch for text retrieval. We first introduce a text encoder (base size) enhanced with RoPE and unpadding, pre-trained in a native 8192-token context (longer than 512 of previous multilingual encoders). Then we construct a hybrid TRM and a cross-encoder reranker by contrastive learning. Evaluations show that our text encoder outperforms the same-sized previous state-of-the-art XLM-R. Meanwhile, our TRM and reranker match the performance of large-sized state-of-the-art BGE-M3 models and achieve better results on long-context retrieval benchmarks. Further analysis demonstrate that our proposed models exhibit higher efficiency during both training and inference. We believe their efficiency and effectiveness could benefit various researches and industrial applications.
@@
Xin Zhang1,2, Yanzhao Zhang1, Dingkun Long1, Wen Xie1, Ziqi Dai1, Jialong Tang1, Huan Lin1 Baosong Yang1, Pengjun Xie1, Fei Huang1, Meishan Zhang, Wenjie Li2, Min Zhang 1Alibaba Group, 2Department of Computing, The Hong Kong Polytechnic University {linzhang.zx,zhangyanzhao.zyz,dingkun.ldk}@alibaba-inc.com https://hf.co/Alibaba-NLP/gte-multilingual-base
@@ 1 Introduction
Concretely, we first introduce a text encoder enhanced with RoPE Su et al. (2024) and unpadding Portes et al. (2023), pre-trained by masked language modeling (MLM) Devlin et al. (2019) via a two-stage curriculum for the native 8,192 tokens context. Based on our encoder, we propose a hybrid TRM capable of generating both elastic dense Kusupati et al. (2022) and sparse vectors for efficient first-stage retrieval, as well as a cross-encoder reranker. We construct them via contrastive learning Wang et al. (2022); Li et al. (2023) with large-scale meticulously curated datasets, providing robust off-the-shelf retrieval models.
We conduct extensive experiments to verify our method. For the text encoder, we evaluate on two natural language understanding (NLU) benchmarks, i.e., XTREME-R Ruder et al. (2021) and GLUE Wang et al. (2018), and show that our encoder outperforms the same-sized previous state-of-the-art XLM-R. For the TRM and reranker, we evaluate on multiple retrieval benchmarks with multilingual and long-context settings, e.g., MIRACL Zhang et al. (2023b) and MLDR Chen et al. (2024), where our models match the performance of state-of-the-art BGE-M3 Chen et al. (2024) and achieve better long-context performance by a smaller size. We will open-source our models and code to facilitate further research and applications.
@@ 2.1 Text Encoder
To construct powerful long-context multilingual text encoder models, we implement several enhancements to BERT Devlin et al. (2019) architecture and train it from scratch using the vocabulary of XLM-R111 https://hf.co/FacebookAI/xlm-roberta-base Conneau et al. (2020) series.
Specifically, we replace the absolute positional embeddings with Rotary Position Embedding (RoPE, Su et al., 2024), and upgrade the FFN to gated linear unit (GLU) Shazeer (2020). To ensure compatibility with libraries like FlashAttention (Dao, 2023), we remove the dropout applied to attention scores. In addition, we pad the token embedding size to be a multiple of 64, which could speedup the model throughput Portes et al. (2023).
@@
Inspired by Portes et al. (2023), we unpad the input batch to reduce redundant computations associated with padding tokens (Figure 2). We use xFormers (Lefaudeux et al., 2022) to implement the variable length attention. It dispatch the attention forward and backward to different kernels222 We adopt the memory-efficient attention Rabe and Staats (2021) in this work. based on the numerical precision, attention head size and device type. We unpad the MLM labels as well to reduce the computation cost of predicting non-masked tokens.
@@
We assemble our multilingual pre-training data from a combination of the following sources: C4 (Raffel et al., 2020), Skypile (Wei et al., 2023) (2021-2023 subsets), mC4 (Xue et al., 2021), CulturaX (Nguyen et al., 2023), Wikipedia Foundation and books (proprietary). We filter them and curate a dataset covering 75 Languages. Appendix Table 7 presents the statistics of our dataset.
@@
We pre-train the model via masked language modeling (MLM) (Devlin et al., 2019)333 We remove the next sentence prediction objective of BERT following (Liu et al., 2019). . The MLM probability is set to $30\%$ Portes et al. (2023). Following Conneau and Lample (2019) and Conneau et al. (2020), the data from different languages is sampled by a multinomial distribution with probabilities $\{q_{i}\}_{i=1\dots N}$, where
@@
Following Portes et al. (2023), we use the learning rate decoupled444 However, Xie et al. (2024) state that the decoupled weight decay is not ideal. We recommend to keep the default setting. AdamW (Loshchilov and Hutter, 2018) with weight decay $51e-$ 5. We disable gradient clipping (set to 0) (Liu et al., 2019). All models are trained on A100 GPU servers by BF16 PyTorch native automatic mixed precision via transformers Wolf et al. (2020). We list the detailed hyper-parameters of each training stage in Appendix A.2 and Table 8. We denote the resulting models as mGTE-MLM-2048/8192.
@@ 2.2 Text Representation Model
We take the encoder output hidden state of the [CLS] token as the dense representation (i.e., embedding) and compute the relevance score by cosine similarity. Our pre-training data (Appendix Table 9) comprise naturally occurring text pairs (e.g., question-answer pairs from Quora and StackExchange, title-content pairs of CommonCrawl), translation pairs Team et al. (2024), and crosslingual instruction tuning data Muennighoff et al. (2022). We train the model with a batch size of $16,384$ and a learning rate of $45e-$ 4 for $240k$ steps. Each batch is sampled from a single data source by the same distribution of Eq.1. The queries (resp. documents) are truncated to the max tokens of $512$ (resp. $1024$). We reverse scale the RoPE base from $160,000$ to $20,000$ to fit the $1024$ context length and acquire the long-context retrieval ability (denotes revNTK, ablation in §3.4). We set $\tau$ of InfoNCE to $0.01$ and only use in-batch negatives. More details refer to Appendix B.3. We denote this contrastive pre-trained model as mGTE-CPT, which is actually an unsupervised embedding model.
@@
Many of recently released models and APIs offer elastic embeddings by Matryoshka representation learning (MRL) Kusupati et al. (2022), providing competitive sub-vectors of embeddings to save index storage and speedup search. Let $\bm{e}\in\mathbb{R}^{H}$ denotes an embedding and $\bm{e}_{:d}$ is the sliced sub-vector from dimension 00 to $d<H$. MRL555 Here we mean the MRL-E in Kusupati et al. (2022). optimizes the weighted sum of multiple losses from different $d$ dimensional sub-vectors, i.e., compute InfoNCE by $s_{d}(\bm{e}^{q}_{:d},\bm{e}^{d}_{:d})$. We add this objective to our TRM fine-tuning stage.
@@ 2.3 Text Reranking Model
We also build a reranker using the cross-encoder architecture. It takes the query and document as input: [CLS] $q$ [SEP] $d$, and directly predicts their relevance score by the [CLS] output state: $s_{\text{rerank}}=\bm{W}\bm{h}_{\text{[CLS]}}$. $\bm{W}\in\mathbb{R}^{H\times 1}$ is randomly initialized. The model is fine-tuned by InfoNCE in one step666 We found that the contrastive pre-training of reranker does not improve the performance. based on our text encoder. Unless otherwise specified, we employ identical data and training settings as the TRM fine-tuning (§2.2). The difference lies in our adjustment of the hard-negatives. We describe the detailed settings in Appendix B.4. We denote this model as mGTE-reranker.
3 Evaluation
3.1 Natural Language Understanding
We evaluate the encoder on the cross-lingual natural language understanding (NLU) benchmark XTREME-R777 We use XTREME-R Ruder et al. (2021) instead of XTREME Hu et al. (2020) since we found the retrieval tasks of XTREME is unstable and difficult to evaluate. Ruder et al. (2021) and the English NLU benchmark GLUE Wang et al. (2018). Results show that our encoder outperforms the same-sized previous state-of-the-art XLM-R Conneau et al. (2020) on all benchmarks.
XTREME-R
GLUE
@@ 3.1 Natural Language Understanding
Table 3: Embedding model performance on MTEB English (Muennighoff et al., 2023), Chinese (Xiao et al., 2023), French Ciancone et al. (2024) and Polish (Poświata et al., 2024). The scores of other models are retrieved from the MTEB online leaderboard. To be consistent with the setting in contrastive pre-training, in retrieval tasks, the max sequence length of the document side is set to 1024. Denote our runs.
@@ 3.2 Text Embedding
Our contrastive pre-training actually yields a text embedding model. To understand the pre-training and fine-tuning of TRM, and to compare with other models, we first run the most popular text embedding benchmark MTEB Muennighoff et al. (2023) as well as its Chinese, French and Polish versions.
@@ B.1 Pre-Training Data
Following previous studies, we create large-scale weakly correlated text pairs from diverse sources. The data are primarily consisted of four parts: English pairs Wang et al. (2022); Li et al. (2023), Chinese pairs Li et al. (2023); Xiao et al. (2023), multilingual pairs (cc-news888 commoncrawl.org/blog/news-dataset-available ), and crosslingual instruction and translation pairs Muennighoff et al. (2022); Team et al. (2024). We filter the data by removing duplicates and low-quality pairs, resulting in a total of 2,938.8M pairs. Table 9 lists the statistics of our contrastive pre-training data (cc-news is separately presented by languages in Table 10).
@@ B.2 Fine-Tuning Data
We collect publicly available high-quality dataset as our fine-tune data as detailed in Table 11. For English, we utilize seven datasets:
  • MS MARCO Nguyen et al. (2016),
  • Natural Questions (NQ) Kwiatkowski et al. (2019),
  • TriviaQA Joshi et al. (2017),
  • HotpotQA Yang et al. (2018),
  • SQuAD Rajpurkar et al. (2016),
  • FEVER Thorne et al. (2018),
  • AllNLI from SimCSE Gao et al. (2021b).
For Chinese, we compile six datasets:
Additionally, we incorporate three multilingual datasets: Mr.TyDi Zhang et al. (2021), MIRACL Zhang et al. (2023b), and MLDR Chen et al. (2024). We exclusively use the trainset of each dataset and employ our contrastive pre-trained model to mine hard negatives.
@@ Appendix C NLU Evaluation
We evaluate our text encoder as well as baselines on the multilingual XTREME-R Ruder et al. (2021) and English GLUE Wang et al. (2018) benchmarks. We describe the fine-tuning setup and the evaluation details in the following subsections. The evaluation scripts are available in our github repo101010 github.com/izhx/nlu-evals .
@@ C.1 XTREME-R
We only run XTREME-R Ruder et al. (2021) in the zero-shot cross-lingual transfer learning setting, where models are fine-tuned on English trainset and tested on multi- and cross-lingual data. We compare our encoder with mBERT-base-cased111111 hf.co/google-bert/bert-base-multilingual-cased and XLM-RoBERTa-base121212 hf.co/FacebookAI/xlm-roberta-base . All models are fine-tuned in the same setting and hyper-parameters.
@@
As XTREME-R has no final release, we implement the evaluation code based on the code of XTREME131313github.com/google-research/xtreme. However, there are some differences in the retrieval evaluation, where our code will deduplicate the retrieval corpus. In addition, we implement the XCOPA in multiple choice, which might be different from XTREME-R. In fine-tuning, if not specified, we use the epoch number of 3, learning rate of 2e-5, batch size of 32, and max sequence length of 128 Hu et al. (2020).
@@
We fine-tune the model on MNLI141414 hf.co/datasets/nyu-mll/glue MNLI subset. Williams et al. (2018) trainset and then evaluate the checkpoint on XNLI151515 hf.co/datasets/facebook/xnli Conneau et al. (2018).
@@
We run this data as the multiple choice task. The model is first trained on SIQA161616 hf.co/datasets/allenai/social_i_qa citesap-etal-2019-social and then COPA171717 hf.co/datasets/aps/super_glue copa split. Roemmele et al. (2011) for 5 epochs on each dataset. The checkpoint of COPA is evaluated on XCOPA181818 hf.co/datasets/cambridgeltl/xcopa Ponti et al. (2020).
@@
We fine-tune the model on the trainset of English by 10 epochs and evaluate on selected WikiANN Rahimi et al. (2019) testsets191919 hf.co/datasets/unimelb-nlp/wikiann .
@@
We fine-tune on the trainset of SQuAD Rajpurkar et al. (2016) v1.1202020 hf.co/datasets/rajpurkar/squad for 3 epochs with the learning rate 3e-5 and max length 384. Then we evaluate the checkpoint on XQuAD212121 hf.co/datasets/google/xquad Artetxe et al. (2020).
@@
We directly evaluate the same checkpoint of XQuAD on MLQA222222 hf.co/datasets/facebook/mlqa Lewis et al. (2020) with the same setting.
@@
We train the model on TyDiQA-GoldP232323 hf.co/datasets/juletxara/tydiqa_xtreme Clark et al. (2020) trainset in the same setting as XQuAD. Then we evaluate the checkpoint on the testset.
@@
We generate the data following their github242424 https://github.com/google-research/google-research/blob/master/dense_representations_for_entity_retrieval/mel/mewsli-x.md#getting-started . This is a updated version so that we can not compare with the results in the XTREME-R paper. We train the model on the English wikipedia (mention, entity)-pairs for 2 epochs with the batch size 64 and max length 64. Then we evaluate the checkpoint in the language agnostic retrieval setting, refer to Ruder et al. (2021) for more details.
@@
This task is actually conducted on XQuAD-R252525 hf.co/datasets/google-research-datasets/xquad_r Roy et al. (2020). We fine-tune the model on the trainset of SQuAD v1.1 in dual-encoder architecture ([CLS] as the embedding) and retrieval setting for 3 epochs with the batch size 16, max query length 96, and max document length 256. Then we evaluate the checkpoint on XQuAD-R in same setting.
@@
We directly evaluate the checkpoint from LAReQA on Tatoeba262626 hf.co/datasets/mteb/tatoeba-bitext-mining Facebook (2019) in the same setting.

### v2

@@ Abstract
We present systematic efforts in building long-context multilingual text representation model (TRM) and reranker from scratch for text retrieval. We first introduce a text encoder (base size) enhanced with RoPE and unpadding, pre-trained in a native 8192-token context (longer than 512 of previous multilingual encoders). Then we construct a hybrid TRM and a cross-encoder reranker by contrastive learning. Evaluations show that our text encoder outperforms the same-sized previous state-of-the-art XLM-R. Meanwhile, our TRM and reranker match the performance of large-sized state-of-the-art BGE-M3 models and achieve better results on long-context retrieval benchmarks. Further analysis demonstrate that our proposed models exhibit higher efficiency during both training and inference. We believe their efficiency and effectiveness could benefit various researches and industrial applications.111 Models are released at https://hf.co/Alibaba-NLP/gte-multilingual-base.
@@
Xin Zhang1,2, Yanzhao Zhang1, Dingkun Long1, Wen Xie1, Ziqi Dai1, Jialong Tang1, Huan Lin1 Baosong Yang1, Pengjun Xie1, Fei Huang1, Meishan Zhangthanks: Corresponding Author, Wenjie Li2, Min Zhang 1Tongyi Lab, Alibaba Group, 2The Hong Kong Polytechnic University {linzhang.zx,zhangyanzhao.zyz,dingkun.ldk}@alibaba-inc.com
@@ 1 Introduction
Concretely, we first introduce a text encoder enhanced with Rotary Position Embedding (RoPE, Su et al., 2024) and unpadding Portes et al. (2023), pre-trained by masked language modeling (MLM) Devlin et al. (2019) via a two-stage curriculum for the native 8,192 tokens context. Based on our encoder, we propose a hybrid TRM capable of generating both elastic dense Kusupati et al. (2022) and sparse vectors for efficient first-stage retrieval, as well as a cross-encoder reranker. We construct them via the contrastive learning objective Wang et al. (2022); Li et al. (2023) with large-scale meticulously curated datasets, providing robust off-the-shelf retrieval models.
We conduct extensive experiments to verify our method. For the text encoder, we evaluate on two natural language understanding (NLU) benchmarks, i.e., XTREME-R Ruder et al. (2021) and GLUE Wang et al. (2018), and show that our encoder outperforms the same-sized previous state-of-the-art XLM-R. For the TRM and reranker, we evaluate on multiple retrieval benchmarks with multilingual and long-context settings, e.g., MIRACL Zhang et al. (2023b) and MLDR Chen et al. (2024), where our models match the performance of state-of-the-art BGE-M3 Chen et al. (2024) and achieve better long-context performance by a smaller size. We open-source our models and code to facilitate further research and applications.
@@ 2.1 Text Encoder
To construct powerful long-context multilingual text encoder models, we implement several enhancements to BERT Devlin et al. (2019) architecture and train it from scratch using the vocabulary of XLM-R222 https://hf.co/FacebookAI/xlm-roberta-base Conneau et al. (2020) series.
Specifically, we replace the absolute positional embeddings with RoPE Su et al. (2024), and upgrade the feedforward network (FFN) to gated linear unit (GLU) Shazeer (2020). To ensure compatibility with libraries like FlashAttention (Dao, 2023), we remove the dropout applied to attention scores. In addition, we pad the token embedding size to be a multiple of 64, which could speedup the model throughput Portes et al. (2023).
@@
Inspired by Portes et al. (2023), we unpad the input batch to reduce redundant computations associated with padding tokens (Figure 2). We use xFormers (Lefaudeux et al., 2022) to implement the variable length attention. It dispatch the attention forward and backward to different kernels333 We adopt the memory-efficient attention Rabe and Staats (2021) in this work. based on the numerical precision, attention head size and device type. We unpad the MLM labels as well to reduce the computation cost of predicting non-masked tokens.
@@
We assemble our multilingual pre-training data from a combination of the following sources: C4 (Raffel et al., 2020), Skypile (Wei et al., 2023) (2021-2023 subsets), mC4 (Xue et al., 2021), CulturaX (Nguyen et al., 2024), Wikipedia Foundation and books (proprietary). We filter them and curate a dataset covering 75 Languages. Appendix Table 7 presents the statistics of our dataset.
@@
We pre-train the model via masked language modeling (MLM) (Devlin et al., 2019)444 We remove the next sentence prediction objective of BERT following (Liu et al., 2019). . The MLM probability is set to $30\%$ Portes et al. (2023). Following Conneau and Lample (2019) and Conneau et al. (2020), the data from different languages is sampled by a multinomial distribution with probabilities $\{q_{i}\}_{i=1\dots N}$, where
@@
Through this method, we could train the model with a large context length in limited resources 555 In our early experiments of English models, we investigated continue training by RetroMAE (Xiao et al., 2022) after MLM-8192. However, we did not observe any improvement. .
@@
Following Portes et al. (2023), we use the learning rate decoupled666 However, Xie et al. (2023b) state that the decoupled weight decay is not ideal. We recommend to keep the default setting. AdamW (Loshchilov and Hutter, 2018) with weight decay $51e-$ 5. We disable gradient clipping (set to 0) (Liu et al., 2019). All models are trained on A100 GPU servers by BF16 PyTorch native automatic mixed precision via transformers Wolf et al. (2020). We list the detailed hyper-parameters of each training stage in Appendix A.2 and Table 8. We denote the resulting models as mGTE-MLM-2048/8192.
@@ 2.2 Text Representation Model
We take the encoder output hidden state of the [CLS] token as the dense representation (i.e., embedding) and compute the relevance score by cosine similarity. Our pre-training data (Appendix Table 9) comprise naturally occurring text pairs (e.g., question-answer pairs from Quora and StackExchange, title-content pairs of CommonCrawl), translation pairs Team et al. (2024), and crosslingual instruction tuning data Muennighoff et al. (2023b). We train the model with a batch size of $16,384$ and a learning rate of $45e-$ 4 for $240k$ steps. Each batch is sampled from a single data source by the same distribution of Eq.1. The queries (resp. documents) are truncated to the max tokens of $512$ (resp. $1024$). We reverse scale the RoPE base from $160,000$ to $20,000$ to fit the $1024$ context length and acquire the long-context retrieval ability (denotes revNTK, ablation in §3.4). We set $\tau$ of InfoNCE to $0.01$ and only use in-batch negatives. More details refer to Appendix B.3. We denote this contrastive pre-trained model as mGTE-CPT, which is actually an unsupervised embedding model.
@@
Many of recently released models and APIs offer elastic embeddings by Matryoshka representation learning (MRL) Kusupati et al. (2022), providing competitive sub-vectors of embeddings to save index storage and speedup search. Let $\bm{e}\in\mathbb{R}^{H}$ denotes an embedding and $\bm{e}_{:d}$ is the sliced sub-vector from dimension 00 to $d<H$. MRL777 Here we mean the MRL-E in Kusupati et al. (2022). optimizes the weighted sum of multiple losses from different $d$ dimensional sub-vectors, i.e., compute InfoNCE by $s_{d}(\bm{e}^{q}_{:d},\bm{e}^{d}_{:d})$. We add this objective to our TRM fine-tuning stage.
@@ 2.3 Text Reranking Model
We also build a reranker using the cross-encoder architecture. It takes the query $q$ and document $d$ together as input: [CLS] $q$ [SEP] $d$, and directly predicts their relevance score by the [CLS] output state: $s_{\text{rerank}}=\bm{W}\bm{h}_{\text{[CLS]}}$. In our experiment, $\bm{W}\in\mathbb{R}^{H\times 1}$ is randomly initialized.
The model is fine-tuned by InfoNCE in one step888 We found that the contrastive pre-training of reranker does not improve the performance. based on our pre-trained 8k-context text encoder model. Unless otherwise specified, we employ identical data and training settings as our TRM fine-tuning stage (§2.2). The difference lies in our adjustment of the hard-negatives. We describe the detailed settings in Appendix B.4. We denote this model as mGTE-reranker.
@@
Table 3: Embedding model performance on MTEB English (Muennighoff et al., 2023a), Chinese (Xiao et al., 2024), French Ciancone et al. (2024) and Polish (Poświata et al., 2024). The scores of other models are retrieved from the MTEB online leaderboard. To be consistent with the setting in contrastive pre-training, in retrieval tasks, the max sequence length of the document side is set to 1024. Denote our runs.
@@
3 Evaluation
3.1 Natural Language Understanding
We evaluate the encoder on the cross-lingual natural language understanding (NLU) benchmark XTREME-R999 We use XTREME-R Ruder et al. (2021) instead of XTREME Hu et al. (2020) since we found the retrieval tasks of XTREME is unstable and difficult to evaluate. Ruder et al. (2021) and the English NLU benchmark GLUE Wang et al. (2018). Results show that our encoder outperforms the same-sized previous state-of-the-art XLM-R Conneau et al. (2020) on all benchmarks.
XTREME-R
GLUE
@@ 3.2 Text Embedding
Our contrastive pre-training actually yields a text embedding model. To understand the pre-training and fine-tuning of TRM, and to compare with other models, we first run the most popular text embedding benchmark MTEB Muennighoff et al. (2023a) as well as its Chinese, French and Polish versions.
@@ 5 Conclusion
Acknowledgements
This work was supported by Alibaba Research Intern Program.
@@ B.1 Pre-Training Data
Following previous studies, we create large-scale weakly correlated text pairs from diverse sources. The data are primarily consisted of four parts: English pairs Wang et al. (2022); Li et al. (2023), Chinese pairs Li et al. (2023); Xiao et al. (2024), multilingual pairs (cc-news101010 commoncrawl.org/blog/news-dataset-available ), and crosslingual instruction and translation pairs Muennighoff et al. (2023b); Team et al. (2024). We filter the data by removing duplicates and low-quality pairs, resulting in a total of 2,938.8M pairs. Table 9 lists the statistics of our contrastive pre-training data (cc-news is separately presented by languages in Table 10).
@@ B.2 Fine-Tuning Data
We collect publicly available high-quality dataset as our fine-tune data as detailed in Table 11. For English, we utilize seven datasets: MS MARCO Nguyen et al. (2016), Natural Questions (NQ) Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017), HotpotQA Yang et al. (2018), SQuAD Rajpurkar et al. (2016), FEVER Thorne et al. (2018), AllNLI from SimCSE Gao et al. (2021b). For Chinese, we compile six datasets: DuReader Qiu et al. (2022), mMARCO-zh Bonifacio et al. (2021), T2-Ranking Xie et al. (2023a), CmedQAv2 Zhang et al. (2018), SimCLUE [^11] https://github.com/CLUEbenchmark/SimCLUE, Multi-CPR Long et al. (2022). Additionally, we incorporate three multilingual datasets: Mr.TyDi Zhang et al. (2021), MIRACL Zhang et al. (2023b), and MLDR Chen et al. (2024). We exclusively use the trainset of each dataset and employ our contrastive pre-trained model to mine hard negatives.
@@ Appendix C NLU Evaluation
We evaluate our text encoder as well as baselines on the multilingual XTREME-R Ruder et al. (2021) and English GLUE Wang et al. (2018) benchmarks. We describe the fine-tuning setup and the evaluation details in the following subsections. The evaluation scripts are available in our github repo121212 github.com/izhx/nlu-evals .
@@ C.1 XTREME-R
We only run XTREME-R Ruder et al. (2021) in the zero-shot cross-lingual transfer learning setting, where models are fine-tuned on English trainset and tested on multi- and cross-lingual data. We compare our encoder with mBERT-base-cased131313 hf.co/google-bert/bert-base-multilingual-cased and XLM-RoBERTa-base141414 hf.co/FacebookAI/xlm-roberta-base . All models are fine-tuned in the same setting and hyper-parameters.
@@
As XTREME-R has no final release, we implement the evaluation code based on the code of XTREME151515github.com/google-research/xtreme. However, there are some differences in the retrieval evaluation, where our code will deduplicate the retrieval corpus. In addition, we implement the XCOPA in multiple choice, which might be different from XTREME-R. In fine-tuning, if not specified, we use the epoch number of 3, learning rate of 2e-5, batch size of 32, and max sequence length of 128 Hu et al. (2020).
@@
We fine-tune the model on MNLI161616 hf.co/datasets/nyu-mll/glue MNLI subset. Williams et al. (2018) trainset and then evaluate the checkpoint on XNLI171717 hf.co/datasets/facebook/xnli Conneau et al. (2018).
@@
We run this data as the multiple choice task. The model is first trained on SIQA181818 hf.co/datasets/allenai/social_i_qa citesap-etal-2019-social and then COPA191919 hf.co/datasets/aps/super_glue copa split. Roemmele et al. (2011) for 5 epochs on each dataset. The checkpoint of COPA is evaluated on XCOPA202020 hf.co/datasets/cambridgeltl/xcopa Ponti et al. (2020).
@@
We fine-tune the model on the trainset of English by 10 epochs and evaluate on selected WikiANN Rahimi et al. (2019) testsets212121 hf.co/datasets/unimelb-nlp/wikiann .
@@
We fine-tune on the trainset of SQuAD Rajpurkar et al. (2016) v1.1222222 hf.co/datasets/rajpurkar/squad for 3 epochs with the learning rate 3e-5 and max length 384. Then we evaluate the checkpoint on XQuAD232323 hf.co/datasets/google/xquad Artetxe et al. (2020).
@@
We directly evaluate the same checkpoint of XQuAD on MLQA242424 hf.co/datasets/facebook/mlqa Lewis et al. (2020) with the same setting.
@@
We train the model on TyDiQA-GoldP252525 hf.co/datasets/juletxara/tydiqa_xtreme Clark et al. (2020) trainset in the same setting as XQuAD. Then we evaluate the checkpoint on the testset.
@@
We generate the data following their github262626 https://github.com/google-research/google-research/blob/master/dense_representations_for_entity_retrieval/mel/mewsli-x.md#getting-started . This is a updated version so that we can not compare with the results in the XTREME-R paper. We train the model on the English wikipedia (mention, entity)-pairs for 2 epochs with the batch size 64 and max length 64. Then we evaluate the checkpoint in the language agnostic retrieval setting, refer to Ruder et al. (2021) for more details.
@@
This task is actually conducted on XQuAD-R272727 hf.co/datasets/google-research-datasets/xquad_r Roy et al. (2020). We fine-tune the model on the trainset of SQuAD v1.1 in dual-encoder architecture ([CLS] as the embedding) and retrieval setting for 3 epochs with the batch size 16, max query length 96, and max document length 256. Then we evaluate the checkpoint on XQuAD-R in same setting.
@@
We directly evaluate the checkpoint from LAReQA on Tatoeba282828 hf.co/datasets/mteb/tatoeba-bitext-mining Facebook (2019) in the same setting.